# Multi-Layer Transformers Gradient Can be Approximated in Almost Linear Time

### v1

@@ 5.3 Accelerating gradient computation of $W_{i}$ and $W_{V_{i}}$
In Section 5.2, we detailed the fast computation of gradients for intermediate variables $T_{i}(X)$. Given that $W_{i}$ is defined as the product $W_{Q_{i}}W_{K_{i}}^{\top}$ (see Definition 1.2), with $W_{Q_{i}}$ and $W_{K_{i}}$ representing the query and key weight matrices, respectively, the gradients of $W_{i}$ and $W_{V_{i}}$ represent all trainable weight matrices in a transformer layer. Consequently, by determining the gradients for $W_{i}$ and $W_{V_{i}}$ across each layer, we achieve almost linear time gradient back-propagation throughout multi-layer transformer models.
@@
The prior study in [6] demonstrated that the gradient of $W_{i}$ can be computed in almost linear time. We extend their findings by adapting their approach to accommodate any loss function $L(X)$ (as defined in Definition 3.1) and further generalize their results to include the gradient computation for both $W_{i}$ and $W_{V_{i}}$ in each transformer layer (Lemma 5.2 and 5.3).
@@
Let $L(X)$ be defined as Definition 3.1. Let $m$ denote the number of self-attention transformer layers (see Definition 1.3). For any $i\in[m]$, let $W_{i}=W_{Q_{i}}W_{K_{i}}^{\top},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th transformer layer. We show that $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}$ can be approximated in $n^{1+o(1)}$ time, with $1/\operatorname{poly}(n)$ approximation error.
@@
Let $L(X)$ be defined as Definition 3.1. Let $m$ denote the number of self-attention transformer layers (see Definition 1.3). For any $i\in[m]$, let $W_{i}=W_{Q_{i}}W_{K_{i}}^{\top},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th transformer layer. We show that $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}$ can be approximated in $n^{1+o(1)}$ time, with $1/\operatorname{poly}(n)$ approximation error.
@@ 5.4 Accelerating gradient computation for multi-Layer transformers
We derive the closed-form gradient for the non-attention components within a transformer layer, namely the $g_{i}$ function defined in Definition 1.3. With the closed-form gradient of $g_{i}$ established in Lemma H.1, we then demonstrate in Lemma 5.4 that the gradient computation for $g_{i}$ can also be achieved in almost linear time. Given that the number of layers $m$ is much smaller than $n$, we can treat $m$ as a constant. Consequently, with respect to running time, since the computation time for gradients on each layer is $n^{1+o(1)}$, we only need to iteratively repeat this procedure for $m$ time. Therefore, the overall running time for computing gradients across the entire model is $m\cdot n^{1+o(1)}$.
@@
We first prove that our $1/\operatorname{poly}(n)$ approximation error statement holds for a single-layer transformer, as evidenced in Lemma H.3. Subsequently, through mathematical induction and leveraging the results of error propagation over the gradient of $g_{i}$, we can show that the approximation error can be bounded by $1/\operatorname{poly}(n)$ for any $m$-layer transformer (Lemma 5.5), where the number of layers $m$ is considered as a constant.
@@
Let $L(X)$ be defined as Definition 3.1. Let $X$ be defined as Definition 1.2. Suppose we have a $m$-layer transformer (see Definition 1.3). Then, for any $i\in[m]$, we can show that,
  • Part 1: running time. Our algorithm can approximate $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}(X)}$, $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}$, and $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}$ in $n^{1+o(1)}$ time.
  • Part 2: error bound. The approximation of the entire transformer model can be bounded by $1/\operatorname{poly}(n)$. Namely, our algorithm output $\widetilde{g}$ satisfies $\|\widetilde{g}-\frac{\mathrm{d}L(X)}{\mathrm{d}X}\|_{\infty}\leq 1/\operatorname{poly}(n)$.
6 Extension
Here, we explore potential enhancements to our algorithm. In Section 6.1, we show the potential extension of multi-head attention and residual connections. In Section 6.2, we offer possibilities to incorporate causal masked attention in almost linear time. In Section 6.3, we show our algorithm can accelerate the prompt tuning process. In Section 6.4, we give the possible integration with system-level attention acceleration techniques.
6.1 Multi-head attention and residual connections
Multi-head attention and residual connections are important components in attention mechanisms. While these components were not involved in our initial analysis for simplicity, incorporating them into our algorithm is straightforward, as detailed in Sections B.1 and B.2. Our algorithm maintains the capability to compute gradients for multi-layer transformers with multi-head attention and residual connection in almost linear time, suggesting that it can be readily adapted to more practical transformer models.
The detailed analysis of incorporating residual connection with our framework can be found in Section J and Lemma J.3. For the synergy with multi-head attention, we provide comprehensive analysis in Section K and Lemma K.2.
6.2 Causal attention mask
The causal attention mask is critical to prevent transformers from “cheating” during training by ensuring future information is not used. The full-rank characteristic of the causal attention mask poses challenges for low-rank approximations. Nevertheless, we have identified a method to accelerate the computation of causal masked attention by exploiting its inherent properties, as demonstrated in [68], remaining almost linear time complexity. A comprehensive explanation is provided in Section B.3.
More detailed analysis of causal attention can be found in Section I and Lemma I.7 and I.8.
6.3 Prompt tuning
@@
6.4 Synergy with system-level attention acceleration
Many contemporary works focus on system-level acceleration of attention mechanisms, often by leveraging caching and mitigating I/O bottlenecks. Our algorithm has the potential to integrate with such advancements. By combining our theoretical improvements in computation time (from $O(n^{2})$ to $n^{1+o(1)}$) with system-level optimizations, the overall efficiency of attention mechanism computation may increase dramatically. We leave the implementation of our method on GPU as future work since there are several coding challenges. More details can be found in Section B.4.
@@ 7 Conclusion
The attention mechanism in transformer models inherently has quadratic time complexity with respect to the input token length. In this work, we propose a novel Algorithm 1, which can approximately train a multi-layer transformer model in almost linear time, introducing only a small error. Moreover, our algorithm is compatible with any loss function, practical sub-modules (residual connection, casual mask, multi-head attention), and gradient-based algorithm and can be seamlessly integrated with other system-level acceleration techniques. While we lack enterprise-scale computational resources for training large language models, our theoretical findings suggest that we can accelerate the training of LLMs in practice.
@@ Acknowledgement
Roadmap. In Section A, we provide further related works of this paper. In Section B, we provide a detailed discussion about several potential extensions of our framework. In Section C, we introduce basic notations and concepts used in our paper, along with the low-rank approximation technique introduced in [5] and [6]. In Section D, we provide details about how we integrate the gradient of $T_{i}(X)$ into matrix form. In Section E, we explain how to apply the low-rank approximation technique to accelerate the computation for the gradient on $T_{i}(X)$. In Section F, we extend the result of [6] to arbitrary loss functions and accelerate the computation of gradient on $W$ via the low-rank approximation technique. In Section G, we calculate the gradient on $W_{V}$ and accelerate the computation of the gradient on $W_{V}$. In Section H, with the help of math induction, we analyze the time complexity and the approximation error across the entire model. In Section I, we discuss how our framework can expand to an attention mechanism with a causal attention mask. In Section J, we provide details about how to integrate our framework with attention mechanism with the residual connection. In Section K, we argue that, with the addition of multi-head attention, our algorithm can still achieve almost linear time gradient computation.
@@ Appendix A More Related Work
[8] introduced attention mechanisms in NLP, enhancing encoder-decoder architecture with variable-length vectors to improve machine translation. Building on this, [71] developed local and global attention variants, further refining NLP tasks. Attention mechanisms found diverse applications: [125] applied them to image captioning, [116]’s Transformer model revolutionized NLP by capturing word relationships, and [115] incorporated attention into graph neural networks. Recent Large Language Model research has focused extensively on attention computation [24, 5, 136, 16, 79, 12, 55]. Studies by [136, 16, 55] use Locality Sensitive Hashing for attention approximation, with [136] offering efficient dot-product attention. [12] and [5] explore static and dynamic attention calculations, while [79] investigates hyperbolic regression regularization. [24] proposes algorithms for reducing attention matrix dimensionality in LLMs. Attention has also been examined from optimization and convergence perspectives [66, 35, 109, 140], investigating word co-occurrence learning [66], regression problems with exponential activation functions [35], attention mechanism evolution during training [109], and the impact of heavy-tailed noise on stochastic gradient descent [140]. Theoretical explorations of attention variants include quantum attention [37], tensor attention [7, 77], and differentially private attention [76, 36, 74].
@@
Various techniques have been developed for model acceleration. One approach involves modifying model architectures to enable faster inference, such as Mamba [33, 23], Linearizing Transformers [135, 92], PolySketchFormer [56], and the Hopfield Model [43, 42, 119, 129, 45, 120, 50] and so on. Another line of work is to prune the weights in a neural network to reduce running time and memory consumption [47, 30, 59, 124, 10, 11, 86, 112, 14, 40, 41, 51, 28, 29, 103, 69]. In addition, specific techniques have been developed to accelerate LLM generation, including KV-Cache compression [39, 62, 134, 143, 82, 26, 130] and speculative decoding [60, 101, 27].
@@ B.1 Multi-head attention
The multi-head attention mechanism was first introduced by [116]. This innovation allows a token to simultaneously attend to multiple positions within the same layer, thereby enriching the model’s capacity for capturing various dependencies. However, this enhanced capability comes with an increase in the size of the attention matrix $f(X)$ from $1\times n\times n$ to $h\times n\times n$, where $h$ is the number of attention heads. To mitigate the computational burden, each head’s vector is derived by splitting the original vector, reducing the dimensionality of each head to $d_{h}:=d/h$. To summarize, the key distinctions between multi-head and single-head attention are (1) an enlarged attention matrix $f(X)$ and (2) a reduced dimensionality $d_{h}$ within each attention head.
@@ B.4 System-level attention acceleration
The attention computing acceleration involves a two-pronged strategy that leverages both system-level improvements (e.g. Flash Attention [22, 20, 100]) and the theoretical time complexity improvements (e.g. our work and [44]).
@@ Appendix C Preliminary on Gradient Calculation
In Section C.1, we list several useful math facts used in the following sections of this paper. In Section C.2, we provide the close forms of the gradient components. In Section C.3, we introduce some mathematical definitions to facilitate understanding of gradient calculations. In Section C.4, we list some low rank approximation technique introduced in [5] and [6]. In Section C.5, we demonstrate that the entries of matrices defined in Section C.3 are bounded.
@@ C.2 Close form of three gradient components
In this section, we show how to derive the close form for the gradient components within each layer of a multi-layer transformer.
@@ C.3 Basic notations for computing gradients
Definition C.11 (Definition of $p(X)$, Definition C.5 in [6]).
@@ C.4 Low rank representations
Using [5]’s polynomial method techniques, we can obtain the following low-rank representation result.
Lemma C.12 (Low rank representation to $f$, Section 3 of [5], Lemma D.1 of [6]).
@@
Lemma C.14 (Low rank representation to $p_{1}(X)$, Lemma D.4 of [6]).
@@
Lemma C.15 (Low rank representation $p_{2}(X)$, Lemma D.5 of [6]).

### v2

@@ 5.3 Accelerating gradient computation of $W_{i}$ and $W_{V_{i}}$
In Section 5.2, we detailed the fast computation of gradients for intermediate variables $T_{i}(X)$. Let $W_{i}:=W_{Q_{i}}W_{K_{i}}^{\top}$, with $W_{Q_{i}}$ and $W_{K_{i}}$ representing the query and key weight matrices, respectively, the gradients of $W_{i}$ and $W_{V_{i}}$ represent all trainable weight matrices in a transformer layer. Consequently, by determining the gradients for $W_{i}$ and $W_{V_{i}}$ across each layer, we achieve almost linear time gradient back-propagation throughout multi-layer transformer models.
@@
The prior study in AS24a [6] demonstrated that the gradient of $W_{i}$ can be computed in almost linear time. We extend their findings by adapting their approach to accommodate general loss function $L(X)$ (as defined in Definition 3.1) and further generalize their results to include the gradient computation for both $W_{i}$ and $W_{V_{i}}$ in each transformer layer (Lemma 5.2 and 5.3).
@@
Let $L(X)$ be defined as Definition 3.1, and $m$ be the number of self-attention transformer layers (Definition 1.3). For any $i\in[m]$, let $W_{i}=W_{Q_{i}}W_{K_{i}}^{\top},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th transformer layer. We show that $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}$ can be approximated in $n^{1+o(1)}$ time, with $1/\operatorname{poly}(n)$ approximation error.
@@
Let $L(X)$ be defined as Definition 3.1, and $m$ be the number of self-attention transformer layers (Definition 1.3). For any $i\in[m]$, let $W_{i}=W_{Q_{i}}W_{K_{i}}^{\top},W_{V_{i}}\in\mathbb{R}^{d\times d}$ denote the attention weight in the $i$-th transformer layer. We show that $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}$ can be approximated in $n^{1+o(1)}$ time, with $1/\operatorname{poly}(n)$ approximation error.
@@ 5.4 Accelerating gradient computation for multi-Layer transformers
We derive the closed-form gradient for the non-attention components within a transformer layer, namely the $g_{i}$ function defined in Definition 1.3. With the closed-form gradient of $g_{i}$ established in Lemma H.1, we then demonstrate in Lemma 5.4 that the gradient computation for $g_{i}$ can also be achieved in almost linear time. Given that the number of layers $m=n^{o(1)}$ is much smaller than $n$ and the computation time for gradients on each layer is $n^{1+o(1)}$, we only need to iteratively repeat this procedure for $m$ times. Therefore, the overall running time for computing gradients across the entire model is $m\cdot n^{1+o(1)}=n^{1+o(1)}$.
@@
We first prove that our $1/\operatorname{poly}(n)$ approximation error statement holds for a single-layer transformer, as evidenced in Lemma H.3. Subsequently, through mathematical induction and leveraging the results of error propagation over the gradient of $g_{i}$, we show that the approximation error can be bounded by $1/\operatorname{poly}(n)$ for any $m$-layer transformer (Lemma 5.5), where the number of layers $m$ is considered small.
@@
Let $L(X)$ be defined as Definition 3.1. Let $X$ be defined as Definition 1.2. Suppose we have a $m$-layer transformer (see Definition 1.3). Then, for any $i\in[m]$, we can show that: ($i$) Running time: Our algorithm can approximate $\frac{\mathrm{d}L(X)}{\mathrm{d}T_{i-1}(X)}$, $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{i}}$, and $\frac{\mathrm{d}L(X)}{\mathrm{d}W_{V_{i}}}$ in $n^{1+o(1)}$ time; ($ii$) Error bound: The approximation of the entire transformer model can be bounded by $1/\operatorname{poly}(n)$. Namely, our algorithm output $\widetilde{g}$ satisfies $\|\widetilde{g}-\frac{\mathrm{d}L(X)}{\mathrm{d}X}\|_{\infty}\leq 1/\operatorname{poly}(n)$.
6 Extensions
Multi-head attention and residual connections.
Multi-head attention and residual connections are important components in attention mechanisms. While these components were not involved in our initial analysis for simplicity, incorporating them into our algorithm is straightforward, as detailed in Sections B.1 and B.2. Our algorithm maintains the capability to compute gradients for multi-layer transformers with multi-head attention and residual connection in almost linear time, suggesting that it can be readily adapted to more practical transformer models. The detailed analysis of incorporating residual connection with our framework can be found in Section J and Lemma J.3. For the synergy with multi-head attention, we provide comprehensive analysis in Section K and Lemma K.2.
Causal attention mask.
The causal attention mask is critical to prevent transformers from “cheating” during training by ensuring future information is not used. The full-rank characteristic of the causal attention mask poses challenges for low-rank approximations. Nevertheless, we have identified a method to accelerate the computation of causal masked attention by exploiting its inherent properties, showing almost linear time complexity. A comprehensive explanation is provided in Section B.3. More detailed analysis can be found in Section I and Lemma I.7 and I.8.
Prompt tuning.
@@ 6 Extensions
Synergy with system-level attention acceleration.
Many contemporary works focus on system-level acceleration of attention mechanisms, often by leveraging caching and mitigating I/O bottlenecks. Our algorithm has the potential to integrate with such advancements. By combining our theoretical improvements in computation time (from $O(n^{2})$ to $n^{1+o(1)}$) with system-level optimizations, the overall efficiency of attention mechanism computation may improve further. We leave the implementation of our method on GPU as future work since there are several coding challenges. More details can be found in Section B.4.
@@ 7 Conclusion
The attention mechanism in transformer models has quadratic time complexity with respect to the input token length. In this work, we proposed a novel Algorithm 1, which can approximately train a multi-layer transformer model in almost linear time, introducing only a small error. Importantly, our algorithm is designed to be compatible with general loss functions, practical sub-modules (residual connection, casual mask, multi-head attention), and general gradient-based algorithms. It may be seamlessly integrated with other system-level acceleration techniques. While we lack enterprise-scale computational resources for training large language models to provide empirical support, our theoretical findings suggest that we can accelerate the training of LLMs in practice.
@@ Acknowledgement
Roadmap. In Section A, we provide further related works of this paper. In Section B, we provide a detailed discussion about several potential extensions of our framework.
In Section C, we introduce basic notations and concepts used in our paper, along with the low-rank approximation technique introduced in AS [5] and AS24a [6]. In Section D, we provide details about how we integrate the gradient of $T_{i}(X)$ into matrix form. In Section E, we explain how to apply the low-rank approximation technique to accelerate the computation for the gradient on $T_{i}(X)$. In Section F, we extend the result of AS24a [6] to arbitrary loss functions and accelerate the computation of gradient on $W$ via the low-rank approximation technique. In Section G, we calculate the gradient on $W_{V}$ and accelerate the computation of the gradient on $W_{V}$. In Section H, with the help of math induction, we analyze the time complexity and the approximation error across the entire model. In Section I, we discuss how our framework can expand to an attention mechanism with a causal attention mask. In Section J, we provide details about how to integrate our framework with attention mechanism with the residual connection. In Section K, we argue that, with the addition of multi-head attention, our algorithm can still achieve almost linear time gradient computation.
@@ Appendix A More Related Work
Attention mechanism.
Attention mechanisms, including self-attention and cross-attention, are pivotal techniques employed in state-of-the-art neural networks. Since it was introduced in VSP+ [105], it has gained widespread adoption across various domains. In particular, it is integral to decoder-only LLMs [88] and the Vision Transformer (ViT) architecture [20]. The former has been instrumental in the remarkable success of LLMs, while the latter has significantly advanced the field of computer vision, encompassing applications such as image generation [87, 112, 113], detection [60], segmentation [127], and layout generation [29, 18, 108]. Moreover, attention mechanism can be integrated into multi-modal models [115, 123, 68, 111], math reasoning [54], diffusion models [83, 69, 42, 24, 78, 70], differential privacy [9, 94, 107, 67, 89, 15, 65, 55, 98] and many other techniques [64, 71, 84, 85, 86, 99, 120, 106].
@@
BCB [8] introduced attention mechanisms in NLP, enhancing encoder-decoder architecture with variable-length vectors to improve machine translation. Building on this, LPM [61] developed local and global attention variants, further refining NLP tasks. Recent Large Language Model research has focused extensively on attention computation [22, 5, 122]. Studies by ZHDK [122], CLP+ [13], KKL [45] use Locality Sensitive Hashing for attention approximation, with ZHDK [122] offering efficient dot-product attention. BSZ [10] and AS [5] explore static and dynamic attention calculations, while LSZ [72] investigates hyperbolic regression regularization. DMS [22] proposes algorithms for reducing attention matrix dimensionality in LLMs. Attention has also been examined from optimization and convergence perspectives [53, 30, 101, 125], investigating word co-occurrence learning [53], regression problems with exponential activation functions [30], attention mechanism evolution during training [101], and the impact of heavy-tailed noise on stochastic gradient descent [125]. Theoretical explorations of attention variants include quantum attention [32], tensor attention [7, 68], and differentially private attention [67, 31, 65].
@@
Various techniques have been developed for model acceleration. One approach involves modifying model architectures to enable faster inference, such as Mamba [28], Linearizing Transformers [121], PolySketchFormer [46], and the Hopfield Model [36, 35, 109, 117, 38, 110, 43, 41] and so on. Another line of work is to prune the weights in a neural network to reduce running time and memory consumption [34, 44, 25, 26, 92, 59, 58]. In addition, specific techniques have been developed to accelerate LLM generation [14, 12, 97, 56].
@@ B.1 Multi-head attention
The multi-head attention mechanism was first introduced by VSP+ [105]. This innovation allows a token to simultaneously attend to multiple positions within the same layer, thereby enriching the model’s capacity for capturing various dependencies. However, this enhanced capability comes with an increase in the size of the attention matrix $f(X)$ from $1\times n\times n$ to $h\times n\times n$, where $h$ is the number of attention heads. To mitigate the computational burden, each head’s vector is derived by splitting the original vector, reducing the dimensionality of each head to $d_{h}:=d/h$. To summarize, the key distinctions between multi-head and single-head attention are (1) an enlarged attention matrix $f(X)$ and (2) a reduced dimensionality $d_{h}$ within each attention head.
@@ B.4 System-level attention acceleration
The attention computing acceleration involves a two-pronged strategy that leverages both system-level improvements (e.g. Flash Attention [21, 19, 90]) and the theoretical time complexity improvements (e.g. our work and HJK+ [37]).
@@ Appendix C Preliminary on Gradient Calculation
In Section C.1, we list several useful math facts used in the following sections of this paper. In Section C.2, we provide the close forms of the gradient components. In Section C.3, we introduce some mathematical definitions to facilitate understanding of gradient calculations. In Section C.4, we list some low rank approximation technique introduced in AS [5] and AS24a [6]. In Section C.5, we demonstrate that the entries of matrices defined in Section C.3 are bounded.
@@ C.2 Close form of three gradient components
We first restate the definition of self-attention, where we denote $W:=W_{Q}W_{K}^{\top}\in\mathbb{R}^{d\times d}$ for simplicity.
Definition C.3 (Self-attention module).
Let $X\in\mathbb{R}^{n\times d}$ denote the input sequence, where $n$ is the number of input tokens and $d$ is the hidden dimension size. Let $W_{V}\in\mathbb{R}^{d\times d}$ be the value weight matrix, and let $W:=W_{Q}W_{K}^{\top}\in\mathbb{R}^{d\times d}$ be the key-query weight matrix. The self-attention function $\mathsf{Attn}(X)$ with weights $W,W_{V}$ is:
$\mathsf{Attn}(X)=\mathsf{Softmax}(XWX^{\top}/d)\cdot X\cdot W_{V}.$
where $\mathsf{Softmax}$ is applied to each row of its input matrix. The attention can be re-written as:
$\mathsf{Attn}(X)=f(X)\cdot X\cdot W_{V},$
Note that the gradient of $W_{Q}$ and $W_{K}$ can easily be calculated from the gradient of $W$, i.e.,
$\frac{\mathrm{d}L(X)}{\mathrm{d}W_{Q}}$ $~{}=\frac{\mathrm{d}L(X)}{\mathrm{d}W}\cdot\frac{\mathrm{d}W}{\mathrm{d}W_{Q}}$
$~{}=\frac{\mathrm{d}L(X)}{\mathrm{d}W}\cdot W_{K}$
where the first step follows from the chain rule, and the second step follows from basic calculus.
Then, we show how to derive the close form for the gradient components within each layer of a multi-layer transformer.
@@ C.3 Basic notations for computing gradients
Definition C.12 (Definition of $p(X)$, Definition C.5 in AS24a [6]).
@@ C.4 Low rank representations
Using AS [5]’s polynomial method techniques, we can obtain the following low-rank representation result.
Lemma C.13 (Low rank representation to $f$, Section 3 of AS [5], Lemma D.1 of AS24a [6]).
@@
Lemma C.15 (Low rank representation to $p_{1}(X)$, Lemma D.4 of AS24a [6]).
@@
Lemma C.16 (Low rank representation $p_{2}(X)$, Lemma D.5 of AS24a [6]).